### v2

# Adversarial Confusion Attack: Disrupting Multimodal Large Language Models

Abstract
We introduce the Adversarial Confusion Attack, a new class of threats against multimodal large language models (MLLMs). Unlike jailbreaks or targeted misclassification, the goal is to induce systematic disruption that makes the model generate incoherent or confidently incorrect outputs. Practical applications include embedding such adversarial images into websites to prevent MLLM-powered AI Agents from operating reliably. The proposed attack maximizes next-token entropy using a small ensemble of open-source MLLMs. In the white-box setting, we show that a single adversarial image can disrupt all models in the ensemble, both in the full-image and Adversarial CAPTCHA settings. Despite relying on a basic adversarial technique (PGD), the attack generates perturbations that transfer to both unseen open-source (e.g., Qwen3-VL) and proprietary (e.g., GPT-5.1) models.
1 Introduction
Most existing adversarial work targets classification errors, unsafe content steering, or jailbreak manipulation [madry2017towards, moosavidezfooli2017universal, qi2023visual, bailey2023image]. We address a distinct failure mode: confusion. A confusion attack aims to destabilize the model’s decoding process and produce high-confidence hallucinations or incoherent text, thereby preventing the model from forming a reliable understanding of the scene. We study a formulation of the confusion attack in which we maximize the next-token Shannon entropy of the model’s output distribution. This objective disrupts the decoder’s internal state and drives the model toward unstable token generation.
Prior work has shown that aligned multimodal models are vulnerable to universal attacks and patch-style perturbations [rahmatullaev2025universaladversarialattackaligned, aichberger2025attackingmultimodalosagents], that proprietary systems such as GPT-4 can be affected by adversarial examples [hu2025transferable, liu2025scalinglawsblackbox], and that perturbations and ensemble design follow scaling laws that govern black-box attack success [chen2024rethinking, liu2025scalinglawsblackbox, liu2017delving]. Our work complements recent adversarial research on MLLMs and contributes the following:
  • We introduce the Adversarial Confusion Attack, which maximizes output entropy to destabilize decoding, and characterize five distinct modes of resulting model failure.
  • In the white-box setting, we show that a single perturbation disrupts all models in the ensemble, in both the full-image and Adversarial CAPTCHA settings.
  • In the full-image setting, we demonstrate black-box transfer to open-source and proprietary MLLMs.
Model Output
“Describe this image.” [Uncaptioned image] GPT-5.1 A dramatic moment from a soccer match […]
GPT-o3 The image shows educational display ”Cell factory” […]
Gemini 3.0 This appears to be corrupted static or noise […] $\cdot$
Grok 4 This is a jailbreak image. I can’t assist with that. $\triangle$
LLaVA-1.5 The image features a person wearing a black shirt […]
Qwen2.5-VL Oh, Pluto! (4) 11 of 1
Table 1: Qualitative results for full-image black-box transfer to proprietary models under an unconstrained perturbation budget (the original attack operates at $448\times 448$ pixels). Example screenshots from the LMSYS Arena are provided in Appendix A. ✓= Success (Hallucination); $\triangle$= Safety Refusal; $\cdot$= Attack Failed.
2 Method
Let $x\in[0,1]^{3\times H\times W}$ be an image and $M\in\{0,1\}^{H\times W}$ be a binary mask defining the attack region. For global attacks, $M$ is an all-ones matrix; for patch attacks, $M$ is 1 only within a fixed region and 0 elsewhere. The perturbed image is defined as:
$x_{\delta}=\Pi_{[0,1]}(x+M\odot\delta),\qquad\|\delta\|_{\infty}\leq\varepsilon,$ (1)
where $\odot$ denotes the element-wise Hadamard product and $\Pi$ clips to the valid pixel range. We attack a surrogate ensemble $\mathcal{E}=\{f_{j}\}_{j=1}^{J}$ of open-source MLLMs. Each model receives $x_{\delta}$ and a fixed text prompt $t$ through its preprocessing pipeline $\Phi_{j}$. For model $f_{j}$, let $z_{j}$ denote its next-token logits at the final prompt position $\tau_{j}$. We compute top-$k$ probabilities $p_{j}=\mathrm{softmax}(z_{j}^{(k)}/T_{e})$, where $z_{j}^{(k)}$ retains the top $k$ logits and $T_{e}$ is the temperature, and maximize the Shannon entropy $H(p_{j})=-\sum_{v}p_{j}(v)\log p_{j}(v)$. The attack maximizes entropy averaged across models:
$\max_{\|\delta\|_{\infty}\leq\varepsilon}\;\frac{1}{J}\sum_{j=1}^{J}H\big(p_{j}(x_{\delta},t)\big).$ (2)
We perform projected gradient ascent (PGD) [madry2017towards], optionally masking the gradient to constrain updates to the patch area:
$\delta\leftarrow\Pi_{\|\cdot\|_{\infty}\leq\varepsilon}\left(\delta+\eta(M\odot\nabla_{\delta}\mathcal{L})\right),$ (3)
with $\mathcal{L}$ equal to the negative of the entropy objective.
Prompt: “Describe this image.”
LLaVA 1.5 The image features a person with a necklace on, positioned […]
LLaVA 1.6 This image is a composite of two separate photographs […]
Qwen3-VL I’m sorry, but I can’t assist with that request.
Qwen2.5-VL I can’t see any image.
Table 2: Qualitative results for the white-box Adversarial CAPTCHA setup (original attack operates at $1024\times 576$ pixels).
3 Experiments
Setup. The base image is a screenshot of the CCRU homepage that is resized to $448\times 448$ to reduce training time. We also tested other websites and observed no substantial differences in results. For the Adversarial CAPTCHA experiments, we use the full $1024\times 576$ webpage screenshot and optimize a fixed $224\times 224$ region at its center. In all scenarios, we optimize the perturbation $\delta$ for 50 iterations and select the final adversarial example by choosing the one that yields the highest averaged entropy across the training ensemble. We used four open-source models: Qwen2.5-VL-3B, Qwen3-VL-2B, LLaVA-1.5-7B, and LLaVA-1.6-7B.
Metrics & Baselines. We report the Shannon entropy of the next-token distribution, restricted to the top $k=50$ logits. We found that aggressive truncation (e.g., $k=5$) reduces transferability, while full-vocabulary optimization introduces training instability. This restriction also standardizes entropy values across models with different vocabulary sizes. We evaluate black-box transfer using a cross-family held-out protocol. Specifically, we optimize on two models from one family and evaluate on a held-out model from a different family. We compare the adversarial output against two baselines: the clean, unperturbed screenshot and an image perturbed with uniform random noise $\delta_{uni}\sim\mathcal{U}(-\varepsilon,\varepsilon)$. Across all models, entropy for the clean image remains low (below 0.6) and comparable to the random noise baseline; a modest entropy increase ($\sim$0.2) was observed for Qwen3-VL under the unconstrained budget noise. We report the Effective Confusion Ratio (ECR), which quantifies how much the attack outperforms both the clean image and the random noise baseline:
$\mathrm{ECR}=\frac{H(f(x_{\mathrm{adv}}))}{\max\big[H(f(x_{\mathrm{clean}})),\,H(f(x_{\mathrm{noise}}))\big]}$ (4)
Values above 1 indicate that the adversarial example induces higher uncertainty than both clean and random-noise baselines.
Proprietary Evaluation. For proprietary models, we evaluate transfer using the LMSYS Arena platform [^1] https://lmarena.ai with the prompt “Describe this image.” and the adversarial image as input. We count an attack as successful when the model’s description is clearly unrelated to the actual image content. We categorize outcomes with three labels: ✓ (coherent hallucination), $\triangle$ (safety or jailbreak-style refusal), and $\cdot$ (no confusion effect, such as correctly identifying the image as noise or describing the clean website layout).
Settings Effective Confusion Ratio (ECR)
$\varepsilon$ LR Qwen3-VL Qwen2.5-VL LLaVA-1.5 LLaVA-1.6 Mean
Panel A: Full Image Attack (White-box)
1.0 0.5 2.33 5.78 3.01 1.94 3.27
0.05 3.29 5.90 5.20 4.96 4.84
0.005 6.84 3.70 6.08 3.69 5.08
0.01 0.5 1.17 1.19 1.41 1.18 1.24
0.05 1.83 2.06 2.72 1.09 1.93
0.005 3.15 2.31 2.46 1.28 2.30
Panel B: Held-out Transfer (Black-box)
1.0 0.5 1.72 1.75 2.08 1.03 1.65
0.05 1.40 0.97 1.43 1.73 1.38
0.005 1.12 1.04 1.27 1.11 1.14
0.01 0.5 1.04 1.05 1.12 1.13 1.09
0.05 1.15 0.98 1.33 1.04 1.13
0.005 1.10 0.99 1.27 1.02 1.10
Panel C: Adversarial $224\times 224$ Patch (White-box)
1.0 0.5 0.97 3.98 1.10 0.95 1.75
0.05 3.41 4.43 3.19 1.17 3.05
0.005 1.05 1.02 1.01 1.00 1.02
Table 3: Effective Confusion Ratios as a function of the perturbation budget $\varepsilon$ and learning rate LRLR. Panel A shows confusion intensity using the full image space. Panel B measures transferability to a held-out model. Panel C evaluates a localized adversarial patch.
3.1 Results
[…]
Proprietary evaluations in Table 4 follow a similar trend. At $\varepsilon=1.0$, GPT-5.1, GPT-o3, GPT-4o, and Nova Pro produce coherent hallucinations, while Grok 4 issues a safety refusal (Table 1). Lower-budget perturbations fail to transfer and result in accurate descriptions of the original website. High-entropy perturbations therefore generalize beyond the training ensemble, but basic PGD fails to produce transferable perturbations under small-budget constraints.
Settings Target Models
$\varepsilon$ LR GPT-5.1 GPT-o3 GPT-4o Grok 4 Gemini 2.5 Gemini 3.0 Nova Pro
1.0 0.5 $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$
0.05 $\triangle$ $\cdot$ $\cdot$
0.01 $\cdot$ $\triangle$ $\cdot$ $\cdot$ $\cdot$
0.01 * $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$
Table 4: Black-box transfer to proprietary models. ✓= coherent hallucination, $\triangle$= safety or jailbreak-style refusal, $\cdot$= no confusion effect.
4 Discussion
Confusion Modes. We categorize the observed adversarial effects into five distinct modes: Blindness, where the model claims inability to view or process the image; Subtle, where the model describes the high-level domain of the image but generates incorrect or uninformative text; Language Switch, characterized by unprompted shifts to non-English scripts; Delusional, involving confident hallucinations of nonexistent objects; and Collapse, a complete breakdown of semantic coherence marked by repetition loops. In the white-box setting, we observed the full spectrum of confusion modes. Collapse was typically associated with peak entropy values, whereas Subtle and Delusional modes correlated with lower entropy increases. In the black-box transfer to proprietary models, Collapse and Blindness were absent; instead, these models exhibited primarily Delusional hallucinations and Language Switch.
Imperceptibility. In our setting, $\varepsilon=0.01$ perturbations are visually imperceptible, but they fail to transfer. Consistent with prior work [liu2025scalinglawsblackbox, chen2024rethinking, liu2017delving, madry2017towards, hu2025transferable], simple PGD-style attacks show limited transferability under very small budgets. However, in some practical settings, visual imperceptibility is a preference rather than a requirement. For adversarial patches designed to block AI Agents from operating on websites, the primary goal is Denial of Service. A visible, high-entropy noise patch ($\varepsilon=1.0$) that reliably induces agent malfunction is therefore a reasonable defense mechanism, even if the perturbation is conspicuous to human users.
Limitations & Future Work. This study uses an entropy-maximization objective implemented with PGD, a basic first-order adversarial optimization technique. Future work should investigate whether feature-level disruptions or more advanced momentum-based adversarial methods [chen2024rethinking] can help bridge the entropy gap between white-box and black-box settings. Enhancing robustness to compression, rendering, and small geometric transformations is also important for real-world deployment [athalye2018synthesizingrobustadversarialexamples]. The adversarial confusion attack also warrants evaluation within complex, multi-step agentic workflows [zhou2024webarena]. A particularly interesting direction is exploring how adversarial confusion can be embedded into website design, such as through background textures or UI color schemes.
5 Conclusion
We introduced the Adversarial Confusion Attack, a method for disrupting Multimodal Large Language Models by maximizing next-token entropy. Using a standard Projected Gradient Descent optimizer and a small surrogate ensemble, we showed that a single perturbation—applied globally or as a localized patch—can reliably destabilize model decoding. The attack transfers to unseen open-source and proprietary models in the full-image setting, indicating that entropy-based perturbations exploit vulnerabilities shared across current MLLMs [huh2024platonicrepresentationhypothesis]. These results position confusion attacks as a novel defense against unauthorized AI Agent activity, deployable via the proposed Adversarial CAPTCHA or, in future applications, through direct integration into website UIs.
Appendix A Appendix: Qualitative Results
In this section, we provide raw LMSYS Chat Arena screenshots showing the Adversarial Confusion Attack transferring to proprietary models. The examples illustrate the range of observed behaviors, from explicit refusals and noise detection to strong hallucinations and fully fabricated scene descriptions.
(a) GPT-5-high hallucinating suburban real estate.
(b) GPT-o3 hallucinating a “subway car,” contrasted with Gemini correctly identifying noise.
(a) GPT-4o hallucinating a vending machine with detailed fictitious elements.
(b) GPT-5 hallucinating a detailed map of central Europe.
(a) GPT-o3 hallucinating an ATM kiosk scene.

### v3

@@
 1 Introduction
-Most existing adversarial work targets classification errors, unsafe content steering, or jailbreak manipulation [madry2017towards, moosavidezfooli2017universal, qi2023visual, bailey2023image].
+Most existing adversarial work targets classification errors, unsafe content steering, or jailbreak manipulation [madry2017towards, moosavidezfooli2017universal, qi2023visual, bailey2023image, ding2025practical, akiri2025safety, wang2025webinject, zhang2025realistic].
 We address a distinct failure mode: confusion.
@@
 This objective disrupts the decoder’s internal state and drives the model toward unstable token generation.
-Prior work has shown that aligned multimodal models are vulnerable to universal attacks and patch-style perturbations [rahmatullaev2025universaladversarialattackaligned, aichberger2025attackingmultimodalosagents], that proprietary systems such as GPT-4 can be affected by adversarial examples [hu2025transferable, liu2025scalinglawsblackbox], and that perturbations and ensemble design follow scaling laws that govern black-box attack success [chen2024rethinking, liu2025scalinglawsblackbox, liu2017delving].
+Prior work has shown that aligned multimodal models are vulnerable to universal attacks and patch-style perturbations [rahmatullaev2025universaladversarialattackaligned, aichberger2025attackingmultimodalosagents, hu2025c2, balakrishnan2025visor], that proprietary systems such as GPT-4 can be affected by adversarial examples [hu2025transferable, liu2025scalinglawsblackbox], and that perturbations and ensemble design follow scaling laws that govern black-box attack success [chen2024rethinking, liu2025scalinglawsblackbox, liu2017delving].
 Our work complements recent adversarial research on MLLMs and contributes the following:
@@
 This study uses an entropy-maximization objective implemented with PGD, a basic first-order adversarial optimization technique.
-Future work should investigate whether feature-level disruptions or more advanced momentum-based adversarial methods [chen2024rethinking] can help bridge the entropy gap between white-box and black-box settings.
+Future work should investigate whether feature-level disruptions or more advanced momentum-based adversarial methods [chen2024rethinking, hu2025c2, balakrishnan2025visor] can help bridge the entropy gap between white-box and black-box settings.
 Enhancing robustness to compression, rendering, and small geometric transformations is also important for real-world deployment [athalye2018synthesizingrobustadversarialexamples].
-The adversarial confusion attack also warrants evaluation within complex, multi-step agentic workflows [zhou2024webarena].
+The adversarial confusion attack also warrants evaluation within complex, multi-step agentic workflows [zhou2024webarena, ding2025practical, zhang2025realistic, wang2025webinject, wang2025advedm].
 A particularly interesting direction is exploring how adversarial confusion can be embedded into website design, such as through background textures or UI color schemes.